OpenStudy (anonymous):

if AB=6 and OC= 4 , find OB

OpenStudy (anonymous):

 THIS IS THE GRAPH

OpenStudy (raden):

use the pythagorean theorem : OB^2 = OC^2 + BC^2 with BC=1/2 AB

OpenStudy (anonymous):

thank you

OpenStudy (raden):

yw

Directrix (directrix):